# Diff-RNTraj: A Structure-aware Diffusion Model for Road Network-constrained Trajectory Generation

### v1

@@ Abstract
Trajectory data is essential for various applications as it records the movement of vehicles. However, publicly available trajectory datasets remain limited in scale due to privacy concerns, which hinders the development of trajectory data mining and trajectory-based applications. To address this issue, some methods for generating synthetic trajectories have been proposed to expand the scale of the dataset. However, all existing methods generate trajectories in the geographical coordinate system, which poses two limitations for their utilization in practical applications: 1) the inability to ensure that the generated trajectories are constrained on the road. 2) the lack of road-related information.
In this paper, we propose a new problem to meet the practical application need, i.e., road network-constrained trajectory (RNTraj) generation, which can directly generate trajectories on the road network with road-related information. RNTraj is a hybrid type of data, in which each point is represented by a discrete road segment and a continuous moving rate. To generate RNTraj, we design a diffusion model called Diff-RNTraj. This model can effectively handle the hybrid RNTraj using a continuous diffusion framework by incorporating a pre-training strategy to embed hybrid RNTraj into continuous representations. During the sampling stage, a RNTraj decoder is designed to map the continuous representation generated by the diffusion model back to the hybrid RNTraj format. Furthermore, Diff-RNTraj introduces a novel loss function to enhance the spatial validity of the generated trajectories. Extensive experiments conducted on two real-world trajectory datasets demonstrate the effectiveness of the proposed model.
@@
Diffusion model, trajectory generation, road network
I Introduction
@@
The widespread deployment of GPS devices in intelligent transportation systems (ITS) makes it convenient to record where a vehicle is. In this case, a sequence of ordered geographical coordinates of a vehicle is termed a trajectory. Mining trajectory data can support the development of various downstream applications in ITS, such as user demand prediction [1, 2], vehicle navigation [3], and personalized route recommendation [4]. Despite the pivotal role of trajectory data, publicly available trajectory datasets are usually limited in scale. This is mainly due to privacy policies and public concern, including regulatory restrictions on trajectory data storage and sharing in many countries and areas, and the concern of users on personal information embedded in trajectories being exposed. The available small-scale trajectory datasets hinder the study of trajectory data mining, since the state-of-the-art trajectory data mining solutions, especially deep-learning based solutions, usually require large-scale datasets to achieve optimal results. To obtain large-scale trajectory data, trajectory generation has gained much attention. It aims to analyze trajectories’ underlying human travel patterns and generate simulated trajectories. The generated trajectories can expand the scale of trajectory datasets, thus improving data availability for trajectory data mining, and finally enhancing the performance of various trajectory-based applications.
@@
Existing efforts for trajectory generation can be categorized into rule-based methods [5, 6] and deep learning methods [7, 8, 9]. Rule-based methods simulate human movement under the assumption that moving behavior adheres to predefined rules. However, they struggle to capture the complexity and uncertainty inherent in human travel patterns, limiting their ability to generate realistic trajectories. By comparison, deep learning methods such as variational auto-encoder (VAE)-based methods[10, 11] and generative adversarial network (GAN)-based methods[12, 13] are more flexible, since they are able to represent the complicated distribution of trajectories. However, all these methods represent trajectories by generating sequences of GPS points (i.e., latitude and longitude) in the geographical coordinate system, causing two limitations when the generated trajectories are utilized by downstream applications. Firstly, directly generating GPS points fails to ensure that the trajectories are on the road. As illustrated in Figure 1(a), the green GPS points $p_{1}\sim p_{6}$ produced by current trajectory generation methods cannot accurately replicate the real moving patterns of vehicles on the road, as a valid and logical GPS point in the trajectory ought to remain on the road consistently. Secondly, most downstream trajectory mining tasks, especially road-related applications, require not only GPS information (e.g., latitude and longitude) but also road-related information (e.g., the road segments vehicle running on), yet such road-related information can not be retrieved directly if we directly generate GPS points.
@@
Figure 1: Compared with the two-stage solution and our end-to-end solution. Two-stage solution: first generate GPS points from Gaussian noise, then perform map matching process. Our solution: directly generate trajectories on the road end-to-end.
Drawing on these observations, we propose a novel task called road network-constrained trajectory generation, i.e., directly generating trajectories on the road network in an end-to-end manner, thereby yielding trajectories directly applicable to subsequent applications. As shown in Figure 1(b), our newly proposed trajectory generation task not only ensures the validity of trajectory but also maintains detailed road-related information on vehicles’ moving behavior. Our proposed generation task calls for defining a new term Road Network-constrained Trajectory (RNTraj).
Figure 2: A RNTraj is a point sequence formed by $\langle q_{1},q_{2},q_{3},q_{4}\rangle$. Each RNTraj point $q$ can be determined by its road segment $e$ and moving ratio $r$, e.g., the road segment of RNTraj point $q_{2}$ is $e_{4}$, and the moving ratio is $r_{2}$. For $q_{3}$, its road segment is $e_{4}$, and the moving ratio is $r_{3}$.
A RNTraj $\tau_{RNTraj}=\langle q_{1},q_{2},\cdots,q_{T}\rangle$ is a sequence of $T$ points in the road network coordinate system. Specifically, each RNTraj point $q=(e,r)$ is determined by a road segment $e$ on which the vehicle is traveling, and a moving ratio $r$ which indicates the proportion of a vehicle’s newly traveled distance along the road segment to the initial remaining untraveled length on that segment. Figure 2 shows an example of RNTraj. The road network-constrained trajectory generation task aims to generate trajectories in the form of RNTraj, ensuring that the generated trajectories are always on the roads and contain road-related information.
In this study, we employ the diffusion model [15, 16] for RNTraj generation, since its capability of generating high-quality, realistic data have been proven in multiple domains, including image generation [17, 18] and audio generation [19, 20]. It demonstrates higher reliability and robustness compared to previous generative methods based on VAE and GAN [21]. Despite the promising performance of diffusion model, generating RNTraj with a diffusion model still faces the following unsolved challenges.
@@
To tackle the aforementioned challenges and achieve end-to-end RNTraj generation effectively, we propose a structure-aware diffusion model for RNTraj generation (Diff-RNTraj). Specifically, Diff-RNTraj can handle hybrid data using proficient continuous diffusion models by introducing a novel idea to convert a hybrid RNTraj into a continuous representation. This involves utilizing a pre-trained strategy to first learn the representation of discrete road segments, followed by concatenating the representation with the moving ratio together to obtain a vectorized RNTraj in the continuous space. Then, during the sampling stage, we propose a hybrid RNTraj decoder to map the generated continuous representations back to RNTraj format. Additionally, to ensure the spatial validity of the generated RNTraj, we incorporate a novel spatial validity loss at each step of the diffusion model’s denoising process to guide model training.
@@
  • Extensive experiments on two real-world trajectory datasets show that compared with other baselines, the trajectories produced by our Diff-RNTraj exhibit the highest level of reasonableness across a range of statistical metrics.
II Related Work
II-A Trajectory Generation
@@
II-B Diffusion Model
@@ III Preliminaries
Definition 2 (Road Network).
@@
  • $length$: The total distance of a road segment in meters.
Definition 3 (Road Network-constrained Trajectory).
We define a road network-constrained trajectory, i.e., RNTraj, of length $T$ as $\tau_{RNTraj}$ = $\langle q_{1},q_{2},\cdots,q_{T}\rangle$, where each RNTraj point $q=(e,r)$ is located in the road coordinate system. Here, $e$ denotes the road segment where $q$ is located, and $r\in[0,1]$ is the moving ratio, representing the proportion of newly traveled distance covered along the road segment to the initial remaining untraveled length on that segment. Taking Figure 2 as an example, we illustrate the meaning of the moving ratio $r$. Here, we have a RNTraj form by $\langle q_{1},\cdots,q_{4}\rangle$ and points $q_{2}$ and $q_{3}$ both pass through road $e_{4}$, so their road segment are both $e_{4}$. For $q_{2}$, it is the first point to through road segment $e_{4}$, so its initial remaining untraveled length is equal to the total distance of road segment $e_{4}$. The distance it traveled on this road is from the start point of $e_{4}$ to the location of $q_{2}$, i.e., $d_{1}$. Therefore, the moving ratio of $q_{2}$ is calculated as $r_{2}=\frac{d_{1}}{e_{4}.length}$. For $q_{3}$, it is not the first point to pass through this road. Consequently, its initial remaining untraveled length is the total length of the road segment minus the distance from the beginning of the road segment to its predecessor RNTraj point, i.e., $e_{4}.length-d_{1}$. The newly traveled distance on this road segment is $d_{2}$, and its moving rate is computed as $\frac{d_{2}}{e_{4}.length-d_{1}}$. Therefore, we define the moving ratio $r_{t}$ for the $t$-th point $q_{t}$ in RNTraj as:
[…]
@@
Given a RNTraj $\tau_{RNTraj}=\langle q_{1},\cdots,q_{T}\rangle$, we can easily calculate the precise GPS coordinates of each RNTraj point according to Algorithm 1.
@@
Input: A RNTraj $\tau_{RNTraj}$ = $\langle q_{1},\cdots,q_{T}\rangle$;
@@
Definition 4 (Road Network Constrained Trajectory Generation).
@@ IV-A RNTraj Vectorization
By combining the road segment representation and the moving ratio along the channel dimension, we can represent a RNTraj $\tau_{RNTraj}$ with length $T$ as a tensor $X\in\mathbb{R}^{T\times(D+1)}$, which is referred to as vectorized RNTraj. The first $D$ dimensions of the vectorized RNTraj correspond to road segment representations, denoted as $X_{e}\in\mathbb{R}^{T\times D}$, and the last dimension represents the moving rate $r$. To scale the moving rate within the range $[-1,1]$, we use the transformation $X{[:,D+1]}=2*r-1$.
@@
IV-B Diffusion Model for Vectorized RNTraj Generation
@@
IV-B1 The Forward Diffusion Process
@@
IV-B2 The Reverse Denoising Process
@@
Figure 4: The architecture of the denoising network, which is stacked by $L$ RDCLs.
IV-B3 Denoising Network
@@
Specifically, the framework of the denoising network is shown in Figure 4. At the $n$-th step of the denoising process, we first implement the positional encoding proposed in DiffWare [25] to project the step $n$ into a $F$-demensional vector $PE[n]\in\mathbb{R}^{F}$. To capture the temporal correlation within trajectory data, we stack $L$ Residual Dilation Convolution Layers (RDCL). These layers are grouped into $\frac{L}{m}$ blocks, each consisting of $m$ layers.
@@ IV-D Model Training
Furthermore, in order to consider sequence dependency and enhance the spatial validity of the generated trajectories, we propose a novel spatial validity loss function that explicitly models the connectivity of generated road segments. Based on the estimated $\hat{X}_{0}$, we reconstruct RNTraj following Section IV-C and obtain a generated road segment sequence $\langle\hat{e}_{1},\hat{e}_{2},\cdots,\hat{e}_{T}\rangle$. To ensure that the adjacent road segments are connected in the UTGraph $\mathcal{G}_{\tau}$, we penalize cases where they are not connected, as these two road segments would be unreachable. Hence, our loss function is defined as:
@@ V Experiments
V-A Dataset
In experiments, we use two publicly available trajectory datasets from Porto, Portugal, and Chengdu, China. The Porto dataset [^1] https://www.kaggle.com/competitions/pkdd-15-predict-taxi-service-trajectory-i/data contains trajectory data of 442 taxis from January 2013 to June 2014. The Chengdu dataset [^2] https://outreach.didichuxing.com/ records taxi trajectory data during October 2016. We download the road network from OpenStreetMap [^3] http://www.openstreetmap.org/ using osmnx tools. We remove the stationary points and filter out trajectories that traveled less than 5 minutes or more than 30 minutes. Then, we apply the map matching algorithm [14] to project GPS points onto the road network and obtain the ground truth of RNTraj. Table I provides statistical information of datasets after pre-processing.
V-B Evaluated Metrics

### v2

@@ Abstract
Trajectory data is essential for various applications. However, publicly available trajectory datasets remain limited in scale due to privacy concerns, which hinders the development of trajectory mining and applications. Although some trajectory generation methods have been proposed to expand dataset scale, they generate trajectories in the geographical coordinate system, posing two limitations for practical applications: 1) failing to ensure that the generated trajectories are road-constrained. 2) lacking road-related information. In this paper, we propose a new problem, road network-constrained trajectory (RNTraj) generation, which can directly generate trajectories on the road network with road-related information. Specifically, RNTraj is a hybrid type of data, in which each point is represented by a discrete road segment and a continuous moving rate. To generate RNTraj, we design a diffusion model called Diff-RNTraj, which can effectively handle the hybrid RNTraj using a continuous diffusion framework by incorporating a pre-training strategy to embed hybrid RNTraj into continuous representations. During the sampling stage, a RNTraj decoder is designed to map the continuous representation generated by the diffusion model back to the hybrid RNTraj format. Furthermore, Diff-RNTraj introduces a novel loss function to enhance trajectory’s spatial validity. Extensive experiments conducted on two datasets demonstrate the effectiveness of Diff-RNTraj.
@@
Spatial-temporal data mining, diffusion model, trajectory generation, road network.
1 Introduction
@@
The widespread deployment of GPS devices in intelligent transportation systems (ITS) makes it convenient to record where a vehicle is. In this case, a sequence of ordered geographical coordinates of a vehicle is termed a trajectory. Mining trajectory data can support the development of various downstream applications in ITS, such as user demand prediction [1, 2, 3], vehicle navigation [4], and personalized route recommendation [5, 6, 7]. Despite the pivotal role of trajectory data, publicly available trajectory datasets are usually limited in scale. This is mainly due to privacy policies and public concern, including regulatory restrictions on trajectory data storage and sharing in many countries and areas, and the concern of users on personal location embedded in trajectories being exposed. The available small-scale trajectory datasets hinder the study of trajectory data mining, since the state-of-the-art trajectory data mining solutions, especially deep-learning based solutions, usually require large-scale datasets to achieve optimal results.
To obtain large-scale trajectory data, the study on trajectory generation has gained much attention [8, 9, 10, 11]. It aims to analyze human potential spatial transfer patterns in cities and generate simulated GPS location sequences, which is vital for many trajectory-based applications. For example, traffic simulation systems utilize the generated trajectory to simulate vehicle movement across road networks, thereby calculating high-fidelity traffic conditions that prompt the advancement of navigation systems [8, 12]. Companies such as Uber leverage synthetic trajectories to expand the scale of trajectory datasets and develop larger models to link users and trajectories [13]. Employing generated trajectory data for research also addresses privacy concerns, as it is devoid of personal information [14].
@@
Existing efforts for trajectory generation can be categorized into rule-based methods [15, 16] and deep learning methods [8, 17, 18]. Rule-based methods simulate human movement under the assumption that moving behavior adheres to predefined rules. However, they struggle to capture the complexity and uncertainty inherent in human travel patterns, limiting their ability to generate realistic trajectories. By comparison, deep learning methods such as variational auto-encoder (VAE)-based methods [19, 20] and generative adversarial network (GAN)-based methods[21, 22] are more flexible, since they are able to represent the complicated distribution of trajectories. However, all these methods represent trajectories by generating sequences of GPS points (i.e., latitude and longitude) in the geographical coordinate system, causing two limitations when the generated trajectories are utilized by downstream applications. Firstly, directly generating GPS points fails to ensure that the trajectories are on the road. As illustrated in Figure 1(a), the green GPS points $p_{1}\sim p_{6}$ produced by current trajectory generation methods cannot accurately replicate the real moving patterns of vehicles on the road, as a valid and logical GPS point in the trajectory ought to stay on the road consistently. Secondly, most downstream trajectory mining tasks, especially road-related applications, require not only precise GPS information (e.g., latitude and longitude) but also road-related information (e.g., the road segments on which the vehicle), yet such road-related information can not be retrieved directly from the generated GPS points.
@@
Figure 1: Comparison between the two-stage solution and our end-to-end solution. Two-stage solution: first generate GPS points, then perform map matching process. Our solution: directly generate trajectories on the road end-to-end.
Drawing on these observations, we propose a novel task called road network-constrained trajectory generation, i.e., directly generating GPS trajectories on the road network in an end-to-end manner, thereby yielding trajectories directly applicable to subsequent applications. As shown in Figure 1(b), our newly proposed trajectory generation task not only ensures the validity of trajectory but also maintains detailed road-related information on vehicles’ moving behavior. Our proposed generation task calls for defining a new term Road Network-constrained Trajectory (RNTraj).
Figure 2: A RNTraj is a point sequence formed by $\langle q_{1},q_{2},q_{3},q_{4}\rangle$. Each RNTraj point $q$ can be determined by road segment $e$ and moving ratio $r$, e.g., the road segment of RNTraj point $q_{2}$ is $e_{4}$, and the moving ratio is $r_{2}$. For $q_{3}$, its road segment is $e_{4}$, and the moving ratio is $r_{3}$.
A RNTraj $\tau_{RN}=\langle q_{1},q_{2},\cdots,q_{T}\rangle$ is a sequence of $T$ points in the road network coordinate system. Specifically, each RNTraj point $q=(e,r)$ is determined by a road segment $e$ on which the vehicle is traveling, and a moving ratio $r$ which indicates the proportion of a vehicle’s newly traveled distance along the road segment to the initial remaining untraveled length on that segment. Figure 2 shows an example of RNTraj. The road network-constrained trajectory generation task aims to generate trajectories in the form of RNTraj, ensuring that the generated trajectories are always on the roads and contain road-related information.
In this study, we propose to generate RNTraj based on diffusion model [24, 25], since diffusion model’s capability of generating high-quality, realistic data have been proven in multiple domains, including image generation [26, 27, 28] and audio generation [29, 30]. It demonstrates higher reliability and robustness compared to previous generative methods based on VAE and GAN [31]. Furthermore, the diffusion model generates data from random Gaussian noise, which eliminates the risk of privacy leakage. Despite the promising performance of diffusion model, generating RNTraj with a diffusion model still faces the following unsolved challenges.
@@
To tackle the aforementioned challenges and achieve end-to-end RNTraj generation effectively, we propose a structure-aware diffusion model for RNTraj generation (Diff-RNTraj). Specifically, Diff-RNTraj can handle hybrid data using continuous diffusion models by introducing a novel idea to convert a hybrid RNTraj into a continuous representation. This involves utilizing a pre-trained strategy to first learn the representation of discrete road segments, followed by concatenating the representation with the moving ratio together to obtain a vectorized RNTraj in the continuous space. Then, during the sampling stage, we propose a hybrid RNTraj decoder to map the generated continuous representations back to RNTraj format. Additionally, to ensure the spatial validity of the generated RNTraj, we incorporate a novel spatial validity loss at each step of the diffusion model’s denoising process to guide model training.
@@
  • Extensive experiments on two real-world trajectory datasets show that compared with other baselines, the trajectories produced by our Diff-RNTraj exhibit the highest level of rationality across a range of statistical metrics.
2 Related Work
2.1 Trajectory Generation
@@
There are some works that focus on generating the road-level trajectory [48, 14, 49], which only generates the road segment sequences. Compared to the RNTraj generation task studied in this paper, road-level trajectory generation is coarse-grained. It cannot obtain precise latitude and longitude information for a vehicle, which will bring shortcomings when analyzing the trajectory behavior, e.g., the travel speed and acceleration patterns.
2.2 Diffusion Model
@@ 3 Preliminaries
Definition 2 (Road Network)
@@
  • $length$: The length of a road segment in meters.
Definition 3 (Road Network-constrained Trajectory)
We define a road network-constrained trajectory, i.e., RNTraj, of length $T$ as $\tau_{RN}$ = $\langle q_{1},q_{2},\cdots,q_{T}\rangle$, where each RNTraj point $q=(e,r)$ is located in the road coordinate system. Here, $e$ denotes the road segment where $q$ is located, and $r\in[0,1]$ is the moving ratio, representing the proportion of newly traveled distance covered along the road segment to the initial remaining untraveled length on that segment.
Taking Figure 2 as an example, we illustrate the meaning of the moving ratio $r$. Here, we have a RNTraj form by $\langle q_{1},\cdots,q_{4}\rangle$ and points $q_{2}$ and $q_{3}$ both pass through road $e_{4}$, so their road segment are both $e_{4}$. For $q_{2}$, it is the first point to through road segment $e_{4}$, so its initial remaining untraveled length is equal to the total distance of road segment $e_{4}$. The distance it traveled on this road is from the start point of $e_{4}$ to the location of $q_{2}$, i.e., $d_{1}$. Therefore, the moving ratio of $q_{2}$ is calculated as $r_{2}=\frac{d_{1}}{e_{4}.length}$. For $q_{3}$, it is not the first point to pass through this road. Consequently, its initial remaining untraveled length is the total length of the road segment minus the distance from the beginning of the road segment to its predecessor RNTraj point, i.e., $e_{4}.length-d_{1}$. The newly traveled distance on this road segment is $d_{2}$, and its moving rate is computed as $\frac{d_{2}}{e_{4}.length-d_{1}}$. Therefore, we define the moving ratio $r_{t}$ for the $t$-th point $q_{t}$ in RNTraj as:
$r_{t}=\begin{cases}\frac{dis(q_{t}.e.start,\,q_{t})}{q_{t}.e.length}&\text{if }q_{t}.e\neq q_{t-1}.e\text{ or }t=1\\ \frac{dis(q_{t-1},\,q_{t})}{q_{t}.e.length-dis(q_{t}.e.start,\,q_{t-1})}&\text{otherwise},\end{cases}$
@@
Given a RNTraj $\tau_{RN}=\langle q_{1},\cdots,q_{T}\rangle$, we can easily calculate the precise GPS coordinates of each RNTraj point according to Algorithm 1.
@@
Input: A RNTraj $\tau_{RN}$ = $\langle q_{1},\cdots,q_{T}\rangle$;
@@
Definition 4 (Road Network Constrained Trajectory Generation)
@@ 4.1 RNTraj Vectorization
By combining the road segment representation and the moving ratio along the channel dimension, we can represent a RNTraj $\tau_{RN}$ with length $T$ as a tensor $X\in\mathbb{R}^{T\times(D+1)}$, which is referred to as vectorized RNTraj. The first $D$ dimensions of the vectorized RNTraj correspond to road segment representations, denoted as $X_{e}\in\mathbb{R}^{T\times D}$, and the last dimension represents the moving rate $r$. To scale the moving rate within the range $[-1,1]$, we use the transformation $X{[:,D+1]}=2*r-1$.
@@
4.2 Diffusion Model for Vectorized RNTraj Generation
@@
4.2.1 The Forward Diffusion Process
@@
4.2.2 The Reverse Denoising Process
@@
Figure 4: The architecture of the denoising network, which is stacked by $L$ residual dilation convolution layers (RDCLs).
4.2.3 Denoising Network
@@
Specifically, the framework of the denoising network is shown in Figure 4. At the $n$-th step of the denoising process, we first implement the positional encoding proposed in DiffWare [35] to project the step $n$ into a $F$-demensional vector $PE[n]\in\mathbb{R}^{F}$:
$\begin{split}PE[n]=[&sin(10^{\frac{0*4}{F/2}}n),sin(10^{\frac{1*4}{F/2}}n),\cdots,sin(10^{\frac{(F/2-1)*4}{F/2}}n),\\ &cos(10^{\frac{0*4}{F/2}}n),cos(10^{\frac{1*4}{F/2}}n),\cdots,cos(10^{\frac{(F/2-1)*4}{F/2}}n)],\end{split}$ (11)
and we implement two fully connected layers on the vector. To capture the temporal correlation within trajectory data, we stack $L$ Residual Dilation Convolution Layers (RDCL). These layers are grouped into $\frac{L}{m}$ blocks, each consisting of $m$ layers.
@@ 4.4 Model Training
Furthermore, in order to consider sequence dependency and enhance the spatial validity of the generated trajectories, we propose a novel spatial validity loss function that explicitly models the connectivity of generated road segments. Based on the estimated $\hat{X}_{0}$, we reconstruct RNTraj following Section 4.3 and obtain a generated road segment sequence $\langle\hat{e}_{1},\hat{e}_{2},\cdots,\hat{e}_{T}\rangle$. To ensure that the adjacent road segments are connected in the UTGraph $\mathcal{G}_{\tau}$, we penalize cases where they are not connected, as these two road segments would be unreachable. Hence, our loss function is defined as:
@@ 5 Experiments
5.1 Dataset
In experiments, we use two publicly available trajectory datasets from Porto, Portugal, and Chengdu, China. The Porto dataset [^1] https://www.kaggle.com/competitions/pkdd-15-predict-taxi-service-trajectory-i/data contains trajectory data of 442 taxis from January 2013 to June 2014. The Chengdu dataset [^2] https://outreach.didichuxing.com/ records taxi trajectory data during October 2016. We download the road network from OpenStreetMap [^3] http://www.openstreetmap.org/ using osmnx tools. We remove the stationary points and filter out trajectories that traveled less than 5 minutes or more than 30 minutes. Then, we apply the map matching algorithm [23] to project GPS points onto the road network and obtain the ground truth of RNTraj. Table I provides statistical information of datasets after pre-processing. Since we used datasets with fixed time intervals between adjacent GPS points, and our generated trajectories also obey the corresponding time patterns, which is helpful for further analysis of trajectory behavior, e.g., varying speeds and acceleration patterns.
5.2 Evaluated Metrics